OpenStudy (anonymous):

An auditorium earned $30,000 in sold−out concert ticket sales. Front section tickets cost $75 per seat and back section tickets cost $50 per seat. The number of front section seats is twice the number of back section seats. How many seats are in the front section?

OpenStudy (anonymous):

@iGreen

OpenStudy (texaschic101):

f = front section tickets b = back section tickets 75f + 50b = 30,000 f = 2b now sub 2b in for f in the other equation 75(2b) + 50b = 30,000 -- distribute 150b + 50b = 30,000 200b = 30,000 b = 30,000/200 b = 150 f = 2b f = 2(150) f = 300 check.. 75f + 50b = 30,000 75(300) + 50(150) = 30,000 22,500 + 7500 = 30,000 30,000 = 30,000 (correct) front seat (f) = 300 back seat (b) = 150

OpenStudy (anonymous):

thanks!

OpenStudy (texaschic101):

no problem :)